OpenStudy (anonymous):

Put three different numbers in the circles so that when you add them at the end of each line you get a square number.

OpenStudy (anonymous):



OpenStudy (anonymous):

(For Fun)

OpenStudy (anonymous):

i dont understand how to solve this

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

X+Y = k^2 Y+Z = m^2 Z+X = n^2

ganeshie8 (ganeshie8):

something like that ? :)

OpenStudy (anonymous):

no it is x+y+z=n^2

ganeshie8 (ganeshie8):

uhh, then its very simple right ? 9 = 1+2+6 so put 1, 2, 6 in circles lol

OpenStudy (anonymous):

Can you show one in the way you said @ganeshie8

ganeshie8 (ganeshie8):

let me think a bit, it can get comlicated :|

ganeshie8 (ganeshie8):

cuz we need to satisfy all 3 X+Y = k^2 Y+Z = m^2 Z+X = n^2

OpenStudy (anonymous):

sure it will more challenging.

OpenStudy (blockcolder):

0, 9, and 16.

ganeshie8 (ganeshie8):

u want all 3 different numbers ? else 2, 2, 2 will work

ganeshie8 (ganeshie8):

oh 0,9, 16 is brilliant...

OpenStudy (anonymous):

but it is not like the question .

ganeshie8 (ganeshie8):

wat do u mean

OpenStudy (anonymous):

oh nice great amazing.

ganeshie8 (ganeshie8):

if u dont allow 0, and dont allow repetition... then it gets a bit complex too..

OpenStudy (anonymous):

Yes it is.

ganeshie8 (ganeshie8):

wud be a nice exercise... to try to find them analytically if possible :)

OpenStudy (blockcolder):

Since by @ganeshie8 's equations, there are 3 equations in 3 variables (X,Y,Z), then we can get solutions if we specify values for k, m, and n. For example, if we solve for X, we have: (X+Y)-(Y+Z)=k^2-m^2 X-Z=k^2-m^2 X+Z=n^2 2X=k^2-m^2+n^2 X=(k^2-m^2+n^2)/2 Y and Z should be the same.

OpenStudy (anonymous):

yes @ganeshie8

OpenStudy (blockcolder):

The only problem now is if the problem specified that the numbers in the circles should be integers. In that case, we can choose values of k, m, and n such that either A. all of them are even B. two of them are odd, and the last one is even This will guarantee that X is an integer, in this case. A similar analysis should work for Y and Z.

OpenStudy (anonymous):

A theory .huh!!

ganeshie8 (ganeshie8):

real numbers is simple as can be seen, but finding integers requires analysis...

OpenStudy (anonymous):

Thanks guys. See you again.Now have to close it.